OpenStudy (anonymous):

what two multiplication facts can you use to solve blank divid 9 equals 7?

OpenStudy (anonymous):

7×9=63 and 9×7=63

OpenStudy (anonymous):

blank divid bye 9 equal 7

OpenStudy (anonymous):

63÷9=7 and 63÷7=9